# DiscoNeRF: Class-Agnostic Object Field for 3D Object Discovery

### v2

@@ Abstract
Neural Radiance Fields (NeRFs) have become a powerful tool for modeling 3D scenes from multiple images. However, NeRFs remain difficult to segment into semantically meaningful regions. Previous approaches to 3D segmentation of NeRFs either require user interaction to isolate a single object, or they rely on 2D semantic masks with a limited number of classes for supervision. As a consequence, they generalize poorly to class-agnostic masks automatically generated in real scenes. This is attributable to the ambiguity arising from zero-shot segmentation, yielding inconsistent masks across views.
In contrast, we propose a method that is robust to inconsistent segmentations and successfully decomposes the scene into a set of objects of any class. By introducing a limited number of competing object slots against which masks are matched, a meaningful object representation emerges that best explains the 2D supervision and minimizes an additional regularization term. Our experiments demonstrate the ability of our method to generate 3D panoptic segmentations on complex scenes, and extract high-quality 3D assets from NeRFs that can then be used in virtual 3D environments.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) [29] have been proposed for novel view synthesis from multiple images and have now become a powerful 3D scene reconstruction tool. One limitation, however, is that they represent the scene geometry in terms of a single 3D density field in which all objects are blended. This makes downstream tasks such as editing or object identification particularly challenging. While there has been work on assigning semantic labels to parts of scenes described in this manner [39, 44, 26], they are limited to classes seen during training. Such classes usually include easily identifiable household items in interior settings and vehicles in street scenes for which labeled data is abundant. Unfortunately, this approach does not scale well to in-the-wild scenes with previously unseen objects. Panoptic Lifting [37] and Instance-NeRF [26] consistently miss objects that are not in a predefined set of classes. When given class-agnostic masks as input, we observed that these methods still fail, as they are unable to deal with the inconsistencies that arise from zero-shot segmentation, as illustrated by Fig. 1(a). Some methods attempt to overcome this limitation by directly learning 3D features instead of labels. For example, DFF [21] and LERF [18] achieve this by distilling 2D features from Dino [4] and CLIP [36] respectively. However, the resulting segmentations tend to be fuzzy and are unable to separate different instances of the same object type. In another line of works, Garfield [19] and Contrastive Lift [3] aim to learn a continuous 3D feature field from 2D masks, which must then be clustered as a post-processing step in order to identify objects. This additional step requires additional hyperparameter tuning and tends to produce imprecise segmentations.
@@
In this work, we propose DiscoNeRF, an approach to segmenting 3D NeRFs in-the-wild without a preset number of classes or user-interaction while still producing sharp segmentations, as illustrated by Fig. 1(b). We start from the observation that foundation models such as SAM [20] or Mask-DINO [23] can assign labels to pixels in 2D images of generic scenes containing an arbitrary number of objects at the cost of these labels being potentially inconsistent across different views of the same scene. Our key insight is that we can efficiently constrain the training of an object field to produce a view-consistent 3D segmentation despite the noisiness of the supervisory signal. This enables us to produce sharp segmentations that are consistent across views, unlike those of the original segmentation algorithm, and for objects belonging to arbitrary classes. These images can then be compared to the 2D segmentation masks returned by the foundation model [20].
Fig. 2 depicts our DiscoNeRF architecture. In addition to the density and color that a standard NeRF predicts at each point in a volume, ours predicts a vector of probabilities of belonging to a particular class. These vectors are then used to render images whose pixels are the probabilities to belong to each class. Note that these object classes are not assigned a semantic meaning, they simply correspond to a coordinate number in the probability vector. To enforce consistency across views between probability images and 2D segmentation masks, we introduce a robust loss function that generalizes the traditional Intersection over Union (IoU) measure to floating point values and relies on the Hungarian algorithm to establish correspondences between regions in the probability images and 2D masks in the segmentations. Minimizing this loss function using the appropriate regularization constraints yields the desired result.
Our contribution is therefore an approach to reconstructing 3D scenes in such a way that objects of any class, as opposed to those from a restricted set of classes seen during training, can be properly segmented. This allows a wide range of additional applications of NeRF scenes, and the identified objects can be used for scene editing or provide spatial understanding to an autonomous system. To ensure reproducibility and encourage future works, we will release our implementation and model weights.
Figure 2: DiscoNeRF pipeline. We add to a standard NeRF (green) an object network (violet) that predicts probabilities of belonging to each object. These are used to render 2D probability images, which are compared to the segmentation masks (blue) produced by a foundation model [20] and made as consistent as possible with them. Crucially, this is done without assigning explicit semantic labels, which improves the generalization abilities of our approach.
@@ 2 Related Work
In this section, we first briefly introduce the original NeRF formulation [29] and then discuss the various approaches to semantic segmentation of a NeRF style representation.
2.1 Neural Radiance Fields (NeRFs)
Given a set of posed images of a scene, NeRFs aim to generate new images of that scene from novel poses. This is done by learning a 3D scene representation in terms of a radiance field, which can be used to render the original images. More specifically, a NeRF model assigns a density $\sigma(\mathbf{x})$ and a view-dependent color $\mathbf{c}(\mathbf{x},\mathbf{d})$ to each 3D point $\mathbf{x}$, where $\mathbf{d}$ is the direction from which $\mathbf{x}$ is seen. The density can be interpreted as the differential probability of a ray terminating at an infinitesimal particle at location $\mathbf{x}$. Given a pixel in an image and the line of sight $\mathbf{r}(t)=\mathbf{o}+t{\bf d}$ for $t$ between the near and far bounds $t_{n}$ and $t_{f}$ going through it, the expected color of that pixel is
@@ 2.1 Neural Radiance Fields (NeRFs)
While the original NeRF implementation suffered from drawbacks in speed and accuracy, recent works have significantly mitigated these issues. These developments make NeRFs an attractive representation for 3D segmentation. A promising line of works improved the training speed of NeRFs by replacing the neural scene representation. TensoRF [6] proposes a reprensentation based on tensor-decomposition to compactly encode scene properties. More recently, InstantNGP [31] has introduced multi-resolution hash encoding to accomplish both efficient and high-fidelity scene modeling, providing a flexible and versatile tool for downstream applications. MipNeRF [1] further improves the view synthesis quality of NeRF by representing the scene at a continuously-valued scale and efficiently rendering anti-aliased conical frustums instead of rays. MipNeRF360 [2] and NeRF++[43] propose to use different re-parametrization techniques to better model general and unbounded scenes. These recent developments have made it possible to reconstruct complex scenes with high visual accuracy. However, the reconstructions remain fully fused in a single neural network and thus unusable for many downstream applications, such as 3D asset generation or editing.
2.2 Segmentation of NeRF Models
@@
In a separate line of work, some methods learn a continuous field of 3D features that allows clustering of different objects in the scene [19, 3, 13, 25, 15]. This representation does not define a specific number of objects and is intrinsically continuous, making it a poor fit for segmenting NeRFs into discrete components. Similarly, feature distillation [21] can be used to learn a 3D field that matches the 2D features from various foundation models. This has been used by recent work to perform text-queries [18] or interactive segmentation [8].
Some previous work demonstrated the potential of NeRFs for 3D semantic modeling [14, 42, 37, 26, 39, 10]. Some of these algorithms [39, 26] query the NeRF model on a sparse grid, and then apply a point-cloud segmentation method based on a 3D U-Net [11]. The generality of these approaches is inherently limited by the pre-defined classes they are trained to identify and they generalize poorly to zero-shot classes.
In a different but related line of work, some methods [35, 12, 28, 41, 32] have been proposed to segment 3D gaussians [17] by assigning labels to individual splats. These approaches do not define a 3D field, unlike NeRF segmentation methods, and due to their very different scene representation it is not straightforward to adapt them to segment scenes reconstructed with NeRFs.
3 Approach
In NeRFs, each point in space is assigned a color and a density, as discussed in Sec. 2.1. Our goal is to also assign a label, which should be the same for all points belonging to the same object. In earlier work [37, 26], this has been tried by making the function $f_{\Theta}$ introduced in Sec. 2.1 to output not only a density and a color but also an object label. This additional output can then be supervised using 2D semantic masks extracted from the ground truth images. However, masks of in-the-wild objects from previously unseen classes are often unavailable or inconsistent across views, making the training of the extended $f_{\Theta}$ difficult.
3.1 Loss Function
To train our object network, we need to design a loss function that is minimized when the binary masks generated from probabilities in the vectors $\mathbf{o}$ are consistent with the 2D segmentation masks generated by the foundation model. The main difficulty being that there is a priori no correspondence between object labels and segmentation masks in individual images.
@@ 3.1 Loss Function
Intuitively, the matching $\gamma$ identifies the most-likely object slot for each input mask. These masks sometimes suffer from over-segmentation arising from zero-shot segmentation. This effect is mitigated by our object field $\mathbf{o}$, which computes probabilities with a softmax layer such that different slots are competing for each pixel and two overlapping representations cannot coexist. Consequently, over-segmented masks will tend to be ignored as masks that are present more often in the supervision signal will spontaneously be preferred during optimization. To further improve the consistency of our segmentation, we introduce an additional regularization term in the following section.
@@ 3.2 Field Regularization
As demonstrated in our experiments, minimizing $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{FP}$ alone produces noisy objects fields. This is expected, as the supervision signal can be highly inconsistent across views. To regularize learning, we implement the object field $\mathbf{o}$ as a 3D hash grid of learnable features [31]. To query $\mathbf{o}(\mathbf{x})$, the 3D point $\mathbf{x}$ is first encoded by interpolating the hash grid features, then these features are decoded by a fully-connected network. Unlike in earlier approaches [26, 37], this modeling allows us to regularize the object field by smoothing the learned features. Specifically, we regularize $\mathbf{o}(\mathbf{x})$ by introducing an additional total variation loss $\mathcal{L}_{TV}$ taken to be
@@
where $h_{i}$ is the learned features for the $i$-th voxel in the grid, and $\mathcal{N}$ defines the neighborhood of adjacent voxels in the grid. We found this additional regularization to be essential to handle the ambiguities of zero-shot 2D segmentations, as we show below in an ablation study. In the presence of conflicting 2D masks such as the ones in Fig. 1, a naive implementation will attempt to fit all masks in the training set, resulting in noisy 3D segmentations. By introducing an additional TV regularizer, our method encourages the emergence of a consistent 3D segmentation. Even if the introduction of this additional loss term can in rare cases lead to overly smooth segmentations, we observed that it improved results both quantitatively and qualitatively.
@@ 3.3 Training
Given a set of posed images as input, we first train a NeRF model to reconstruct the scene by sampling random rays across different views. We use recent techniques that help make NeRF reconstruction magnitudes faster and more accurate. We train a proposal network [2] to improve the ray sampling strategy, and employ hash grid features [31] to speed up training. We exploit the fact that a coarse-to-fine activation of hash grids [24] improves the robustness of reconstruction at a minimal cost in training speed.
Then, we freeze the weights of the NeRF model and train the object network. Unlike in the previous step, this optimization is performed by sampling random views and considering all the masks associated to these views jointly. We then minimize $\mathcal{L}_{\gamma}+\lambda_{FP}\mathcal{L}_{FP}+\lambda_{TV}\mathcal{L}_{TV}$, where $\lambda_{FP}$ and $\lambda_{TV}$ are taken to be 0.01 unless explicitly stated otherwise.
@@ 4 Experiments
Our implementation is based on torch-ngp [38], which incorporates the speedup of Instant-NGP [31]. To generate the binary segmentation masks of $M_{m}$ of Sec. 3, we use the Segment Anything Model (SAM) [20]. For object-centric scenes, we generate an additional background mask by extracting the pixels that lie outside of the unit sphere in 3D using the reconstructed NeRF.
@@
Figure 4: Comparing 3D NeRF decomposition models for the scenes of counter and room.
@@ 4.1 Comparative Evaluation
We evaluate our method quantitatively on the Mip-NeRF 360 dataset [2], which features both indoor and outdoor scenes containing a wide variety of common and less common objects. To this end, we render our object field to an image and select the $\operatorname*{argmax}$ per pixel to produce discrete segmentation masks. We then compare them against hand-generated ground-truth masks for each scene except the stump scene, which has no object and was therefore excluded from our experiments. To ensure the reproducibility of our experiments, we will publish these masks and all the results presented along with our implementation.
@@
Metrics.
along with its symmetric counterpart $SBD(\mathcal{P},\mathcal{G})=\min\{BD(\mathcal{P},\mathcal{G}),BD(\mathcal{G},\mathcal{P})\}$. An in-depth analysis of these two metrics can be found in [7].
Baselines.
For comparison purposes, we implement a modernized version of Distilled Feature Fields (DFF) [21], which jointly learns a NeRF model and a 3D field replicating the features produced by a foundation model in a process called feature distillation. Specifically, we implement DFF with distillation of the recent DINOv2 features [33], as opposed to DINO features in the original method [4]. We will refer to this method as DFFv2. To make the distillation tractable, we first reduce the dimensionality of DINOv2 features to 64 with principal component analysis. We then produce 2D segmentations by rendering the distilled features and clustering them.
We also compare our method against Panoptic Lifting [37] and Instance-NeRF [26], two recent methods for semantic segmentation of NeRF models that rely on 2D and 3D segmentation models pretrained on specific object classes. Panoptic Lifting [37] jointly trains a NeRF and a multi-layer perceptron to match 2D semantic and instance maps. Additionally, we modified Panoptic Lifting to use masks produced by SAM. We refer to this method as PL+SAM. Note that the 2D prediction confidence and the semantic losses $\mathcal{L}_{sem}$ and $\mathcal{L}_{con}$ of the original paper cannot be used in this context due to the absence of ground-truth semantic classes. Instance-NeRF [26] fuses inconsistent 2D segmentations by querying a NeRF on a 3D grid and applying a point-cloud segmentation model. Combining Instance-NeRF with SAM does not improve its performance, as it does not only rely on semantic 2D masks but also on a semantic 3D point-cloud segmentation network.
@@
Results.
Similarly, Panoptic Lifting relies on a semantic panoptic segmentation of the scene [9] and fails to properly segment novel objects for in-the-wild scenes. The results of our implementation of PL+SAM, which uses the same masks as our method as supervision, show that Panoptic Lifting is unable to derive a meaningful 3D segmentation from class-agnostic masks. In complex scenes, several objects are wrongfully clustered together or associated with the background. In contrast, our method is fully class-agnostic and can be applied to scenes with zero-shot objects.
@@ Results.
Figure 6: Additional results on the LERF [18] dataset.
@@
Figure 7: Applications. Given some images (left), we perform simple editing operations to synthesize new images (right) by conditioning NeRF rendering with the object field value.

### v3

# Enforcing View-Consistency in Class-Agnostic 3D Segmentation Fields

@@ Abstract
Radiance Fields have become a powerful tool for modeling 3D scenes from multiple images. However, they remain difficult to segment into semantically meaningful regions. Some methods work well using 2D semantic masks, but they generalize poorly to class-agnostic segmentations. More recent methods circumvent this issue by using contrastive learning to optimize a high-dimensional 3D feature field instead. However, recovering a segmentation then requires clustering and fine-tuning the associated hyperparameters. In contrast, we aim to identify the necessary changes in segmentation field methods to directly learn a segmentation field while being robust to inconsistent class-agnostic masks, successfully decomposing the scene into a set of objects of any class.
By introducing an additional spatial regularization term and restricting the field to a limited number of competing object slots against which masks are matched, a meaningful object representation emerges that best explains the 2D supervision. Our experiments demonstrate the ability of our method to generate 3D panoptic segmentations on complex scenes, and extract high-quality 3D assets from radiance fields that can then be used in virtual 3D environments.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) [30] and 3D Gaussian Splatting [18] have been proposed for novel view synthesis from multiple images and have now become powerful 3D scene reconstruction tools. One limitation, however, is that they represent the scene geometry in terms of a single 3D density field in which all objects are blended. This makes downstream tasks such as editing or object identification particularly challenging. While there has been work on assigning semantic labels to parts of scenes described in this manner [41, 46, 27], the majority of early works are limited to classes seen during training. Such classes usually include easily identifiable household items in interior settings and vehicles in street scenes for which labeled data is abundant. Unfortunately, this approach does not scale well to in-the-wild scenes with previously unseen objects. Panoptic Lifting [38] and Instance-NeRF [27] consistently miss objects that are not in a predefined set of classes. When given class-agnostic masks as input, we observed that these methods still fail, as they are unable to deal with the inconsistencies that arise from zero-shot segmentation, as illustrated by Fig. 1(a). More recently, some methods have attempted to overcome this limitation by directly learning 3D features instead of labels. For example, DFF [22] and LERF [19] achieve this by distilling 2D features from Dino [4] and CLIP [37] respectively. However, the resulting segmentations tend to be fuzzy and are unable to separate different instances of the same object type. In another line of works, Garfield [20] and Contrastive Lift [3] aim to learn a continuous 3D feature field from 2D masks, which must then be clustered as a post-processing step in order to produce a segmentation field. This additional step tends to produce imprecise segmentations unless additional hyperparameter are carefully set, such as the desired scale of Garfield [20] or the clustering distance threshold in Contrastive Lift [3].
@@
In this work, we propose DiscoNeRF, an approach to discover 3D objects in an open-world environment. DiscoNeRF segments 3D radiance fields without a preset number of classes or user-interaction while still producing sharp segmentations, as illustrated by Fig. 1(b). We observe that foundation models such as SAM [21] or Mask-DINO [24] can assign labels to pixels in scenes containing an arbitrary number of objects at the cost of these labels being potentially inconsistent across views of the same scene. While operations such as mask dilation or disconnected component removal can sometimes help reduce the signal-to-noise ratio in these 2D masks, they are insufficient for significant artifacts and view-inconsistent representations. Our key insight is that we can efficiently constrain the training of an object field to produce a view-consistent 3D segmentation despite the noisiness of the supervisory signal. This enables us to produce sharp segmentations that are consistent across views, unlike those of the original segmentation algorithm, and for objects belonging to arbitrary classes. These images can then be compared to the 2D segmentation masks returned by the foundation model [21].
Fig. 2 depicts our architecture. In addition to the density and color that a standard radiance field predicts at each point in a volume, ours predicts a vector of probabilities of belonging to a particular object. These vectors are then used to render images whose pixels are the probabilities to belong to each class. Note that these object classes are not assigned a semantic meaning, they simply correspond to a channel in the probability vector. To enforce consistency across views between probability images and 2D segmentation masks, we introduce a robust loss function that generalizes the traditional Intersection over Union (IoU) measure to floating point values and relies on the Hungarian algorithm to establish correspondences between regions in the probability images and 2D masks in the segmentations. Minimizing this loss function using the appropriate regularization constraints yields the desired result.
Our contribution is therefore an approach to reconstructing 3D scenes in such a way that objects of any class can be properly segmented. We achieve this by matching mask with object channels and introducing an additional spatial-consistency regularizer. This allows a wide range of additional applications of reconstructed scenes, and the identified objects can be used for scene editing or provide spatial understanding to an autonomous system.
Figure 2: DiscoNeRF pipeline. We add to a radiance field (green) an object network (violet) that predicts probabilities of belonging to each object. These are used to render 2D probability images, which are compared to the segmentation masks (blue) produced by a foundation model [21]. We introduce three losses $\mathcal{L}_{TV},\mathcal{L}_{\gamma}$ and $\mathcal{L}_{FP}$, designed to be robust to the inconsistency in the supervision signal.
@@ 2 Related Work
In this section, we first briefly introduce the radiance field formulation [30] and then discuss the various approaches to semantic segmentation of such a representation.
2.1 Radiance Fields
Given a set of posed images of a scene, radiance fields [30] aim to generate new images of that scene from novel poses. This is done by learning a 3D scene representation which can be used to render the original images. More recently, 3D Gaussian Splatting [18] has successfully proposed an explicit radiance field representation. In both cases, a radiance field model assigns a density $\sigma(\mathbf{x})$ and a view-dependent color $\mathbf{c}(\mathbf{x},\mathbf{d})$ to each 3D point $\mathbf{x}$, where $\mathbf{d}$ is the direction from which $\mathbf{x}$ is seen. The density can be interpreted as the differential probability of a ray terminating at an infinitesimal particle at location $\mathbf{x}$. Given a pixel in an image and the line of sight $\mathbf{r}(t)=\mathbf{o}+t{\bf d}$ for $t$ between the near and far bounds $t_{n}$ and $t_{f}$ going through it, the expected color of that pixel is
@@ 2.1 Radiance Fields
While the original NeRF implementation suffered from drawbacks in speed and accuracy, recent works have significantly mitigated these issues. A promising line of works improved the training speed of NeRFs by replacing the neural scene representation. TensoRF [6] proposes a reprensentation based on tensor-decomposition to compactly encode scene properties. More recently, InstantNGP [32] has introduced multi-resolution hash encoding to accomplish both efficient and high-fidelity scene modeling, providing a flexible and versatile tool for downstream applications. MipNeRF [1] further improves the view synthesis quality of NeRF by representing the scene at a continuously-valued scale and efficiently rendering anti-aliased conical frustums instead of rays. MipNeRF360 [2] and NeRF++[45] propose to use different re-parametrization techniques to better model general and unbounded scenes. Finally, explicit methods such as SVRaster [39], Radiant Foam [15], or 3D Gaussian Splatting [18] have proposed explicit representations that greatly speed up rendering speed whilst preserving state-of-the-art visual quality. These recent developments have made it possible to reconstruct complex scenes with high visual accuracy. However, the reconstructions remain fully fused in a single reconstruction and thus unusable for many downstream applications, such as 3D asset generation or editing.
2.2 Segmentation fields
@@
In a separate line of work, some methods use contrastive learning to optimize a continuous field of 3D features that allows clustering of different objects in the scene [20, 3, 13, 26, 16]. These representations learn a 3D feature field and necessitate additional clustering to produce a segmentation field, requiring carefully set hyperparameters to produce the desired results. Furthermore, they do not define a specific number of objects and is intrinsically continuous, making it a poor fit for segmenting NeRFs into discrete components. Similarly, feature distillation [22] can be used to learn a 3D field that matches the 2D features from various foundation models. This has been used by recent work to perform text-queries [19] or interactive segmentation [8].
Some previous work demonstrated the potential of radiance fields for 3D semantic modeling [14, 44, 38, 27, 41, 10]. Some of these algorithms [41, 27] query the NeRF model on a sparse grid, and then apply a point-cloud segmentation method based on a 3D U-Net [11]. The generality of these approaches is inherently limited by the pre-defined classes they are trained to identify and they generalize poorly to zero-shot classes. In a different but related line of work, some methods [36, 12, 29, 43, 33] have been proposed to segment 3D gaussians [18] by assigning labels to individual splats. These approaches do not define a 3D field, unlike NeRF segmentation methods, and due to their vastly different scene representation it is not straightforward to adapt them to segment scenes reconstructed with NeRFs.
@@
3 Approach
In radiance fields, each point in space is assigned a color and a density, as discussed in Sec. 2.1. Our goal is to also assign a label, which should be the same for all points belonging to the same object. In earlier work [38, 27], this has been tried by making the function $f_{\Theta}$ introduced in Sec. 2.1 to output not only a density and a color but also an object label, thus implicitly defining a segmentation field. This additional output can then be supervised using 2D semantic masks extracted from the ground truth images. However, masks of in-the-wild objects from previously unseen classes are often unavailable or inconsistent across views, making the training of the extended $f_{\Theta}$ difficult.
3.1 Loss Function
To train our object network, we need to design a loss function that is minimized when the binary masks generated from probabilities in the vectors $\mathbf{o}$ are consistent with the 2D segmentation masks generated by the foundation model. The main difficulty being the noisiness in the supervision signal, and the fact that there is a priori no correspondence between object labels and segmentation masks in individual images.
@@ 3.1 Loss Function
Intuitively, the matching $\gamma$ identifies the most-likely object slot for each input mask. These masks sometimes suffer from over-segmentation arising from zero-shot segmentation. This effect is mitigated by our object field $\mathbf{o}$, which computes probabilities with a softmax layer such that different slots are competing for each pixel and two overlapping representations cannot coexist. Consequently, over-segmented masks will tend to be ignored as masks that are present more often in the supervision signal will spontaneously be preferred during optimization.
While a comparable mask matching is present in [38], it is used on masks with limited classes and fails on class-agnostic segmentations. As demonstrated in our experiments, our proposed $\alpha$, $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{FP}$ are necessary to robustly deal with this inconsistent 2D signal. To further improve the consistency of our segmentation, we introduce an additional regularization term in the following section.
@@ 3.2 Field Regularization
As demonstrated in our experiments, minimizing $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{FP}$ alone can still produce noisy segmentation fields. To regularize learning, we implement the object field $\mathbf{o}$ as a 3D hash grid of learnable features [32]. To query $\mathbf{o}(\mathbf{x})$, the 3D point $\mathbf{x}$ is first encoded by interpolating the hash grid features, then these features are decoded by a fully-connected network. Unlike in earlier approaches [27, 38], this modeling allows us to regularize the object field by smoothing the learned features. Specifically, we regularize $\mathbf{o}(\mathbf{x})$ by introducing an additional total variation loss $\mathcal{L}_{TV}$ taken to be
@@
where $h_{i}$ is the learned features for the $i$-th voxel in the grid, and $\mathcal{N}$ defines the neighborhood of adjacent voxels in the grid. We found this additional regularization to be essential to handle the ambiguities of zero-shot 2D segmentations, as we show below in an ablation study. In the presence of conflicting 2D masks such as the ones in Fig. 1, a naive implementation will attempt to fit all masks in the training set, resulting in noisy 3D segmentations. By introducing an additional TV regularizer, our method encourages the emergence of a consistent 3D segmentation. Even if the introduction of this additional loss term can in rare cases lead to overly smooth segmentations, we observed that it overall improved results both quantitatively and qualitatively.
@@ 3.3 Training
Given a set of posed images as input, we first train a neural radiance field to reconstruct the scene by sampling random rays across different views. We use recent techniques that help make radiance field reconstruction magnitudes faster and more accurate. We train a proposal network [2] to improve the ray sampling strategy, and employ hash features [32] to speed up training. We use a coarse-to-fine activation of hash grids [25], improving the robustness of reconstruction at a minimal cost in training speed.
Then, we freeze the weights of the model and train the object network. Unlike in the previous step, this optimization is performed by sampling random views and considering all the masks associated to these views jointly. We then minimize $\mathcal{L}_{\gamma}+\lambda_{FP}\mathcal{L}_{FP}+\lambda_{TV}\mathcal{L}_{TV}$, where $\lambda_{FP}$ and $\lambda_{TV}$ are taken to be 0.01 unless explicitly stated otherwise.
@@ 4 Experiments
Our implementation is based on torch-ngp [40], which incorporates the speedup of Instant-NGP [32]. To generate the segmentation masks of $M_{m}$ in Sec. 3, we use SAM [21]. For object-centric scenes, we generate a background mask from the pixels that lie outside of the 3D unit sphere.
@@
Figure 4: Comparing 3D radiance field decomposition models for the scenes of counter and room.
@@ 4.1 Comparative Evaluation
We evaluate our method quantitatively on the Mip-NeRF 360 dataset [2], which features both indoor and outdoor scenes containing a wide variety of common and less common objects. To this end, we render our object field to an image and select the $\operatorname*{argmax}$ per pixel to produce discrete segmentation masks. We then compare them against hand-generated ground-truth masks for each scene except the stump scene, which has no object and was therefore excluded from our experiments.
@@
Metrics.
along with its symmetric counterpart $SBD(\mathcal{P},\mathcal{G})=\min\{BD(\mathcal{P},\mathcal{G}),BD(\mathcal{G},\mathcal{P})\}$, as detailed in [7].
Baselines.
For comparison purposes, we implement a modernized version of Distilled Feature Fields (DFF) [22], which jointly learns a NeRF model and a 3D field replicating the features produced by a foundation model in a process called feature distillation. Specifically, we implement DFF with distillation of the DINOv2 features [34], as opposed to DINO features in the original method [4]. We will refer to this method as DFFv2. To make the distillation tractable, we first reduce the dimensionality of DINOv2 features to 64 with PCA, then produce 2D segmentations by rendering the distilled features and clustering them.
We also compare our method against Panoptic Lifting [38] and Instance-NeRF [27], two methods for semantic segmentation of NeRF models that rely on 2D and 3D segmentation models pretrained on specific object classes. Panoptic Lifting [38] jointly trains a NeRF and a multi-layer perceptron to match 2D semantic and instance maps. Additionally, we modified Panoptic Lifting to use masks produced by SAM. We refer to this method as PL+SAM. Note that the 2D prediction confidence and the semantic losses $\mathcal{L}_{sem}$ and $\mathcal{L}_{con}$ of the original paper cannot be used in this context due to the absence of ground-truth semantic classes. Instance-NeRF [27] fuses inconsistent 2D segmentations by querying a NeRF on a 3D grid and applying a point-cloud segmentation model. Combining Instance-NeRF with SAM does not improve its performance, as it does not only rely on semantic 2D masks but also on a semantic 3D point-cloud segmentation network. Note that we do not compare with contrastive methods, as they only learn a feature field and require fine-tuning additional clustering hyperparameters to generate a usable segmentation field.
@@
Figure 6: Additional results on the LERF [19] dataset. In some cases, high ambiguity in the mask generation can still lead to artifacts or fuzzy segmentations, as can be seen at the bottom right of the Waldo kitchen scene.
@@
Figure 7: Applications. Given some images (left), we perform simple editing operations to synthesize new images (right) by conditioning radiance field rendering with the object field value.
Results.
Similarly, Panoptic Lifting relies on a semantic panoptic segmentation of the scene [9] and fails to properly segment novel objects for in-the-wild scenes. The results of our implementation of PL+SAM, which uses the same masks as our method as supervision, show that it is unable to derive a meaningful 3D segmentation from class-agnostic masks. In complex scenes, several objects are wrongfully clustered together or associated with the background.